Bill has a 7-year, 8.9% bond issue selling for $991.78.

  1. What is the yield to maturity on this bond?

  1. Is the bond selling at a premium or a discount?

  1. What is the current yield on this bond?

Remember: bond interest payments are semi-annual, and maturity and face value are $1,000
Answer #1
1) Yield to maturity is 9.06%
Working:
Yield to maturity = =rate(nper,pmt,-pv,fv)*2 Where,
= =rate(N6,N7,-N8,N9)*2 nper 7*2 = 14
= 9.06% pmt 1000*8.9%*6/12 = $       44.50
pv = $     991.78
fv = $       1,000
2) Bond is selling at discount.
Working:
When bond is selling below face value, it is selling at discount.
In the current case, bond's selling price $ 991.78 is below face value of $ 1,000.
So, it is selling at discount.
3) Current Yield 8.97%
Working:
Annual coupon payment in cash = Face Value * Coupon rate
= 1000*8.9%
= $    89.00
Current Yield = Annual coupon paid in cash/Current Price
= $    89.00 / $ 991.78
= 8.97%